Bearclaws72:

Write an expression that is equivalent to three over four (5z+ 16). 15 over fourz + 16 15 over fourz + 12 three over fourz + 16 three over fourz + 12

Vocaloid:

use the distributive property if A(B+C) = A*B + A*C use this pattern to find: (3/4) (5z + 16) = ?

Bearclaws72:

15.75

Vocaloid:

you can't combine like terms (5z+16)

Vocaloid:

(3/4) *(5z) = ?

Bearclaws72:

0.75

Vocaloid:

(3/4)*5 = ?

Bearclaws72:

3.75

Vocaloid:

that's good but we can leave it as a fraction 15/4

Vocaloid:

so (3/4)*5z = (15/4)z what is (3/4)*16 = ?

Bearclaws72:

12

Vocaloid:

Good so your answer is (5/4)z + 12

Bearclaws72:

k